OpenStudy (anonymous):

Prove x^2 (dy/dx) - xy = 1 + cos(y/x) is a homogeneous differential equation.

OpenStudy (anonymous):

do u know whats homogeneous diff equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats ur definition of homogeneous diff equation?

OpenStudy (anonymous):

If \[\frac{ dy }{ dx } = \frac{ M(x,y) }{ N(x,y) }\] then its said to be in homogeneous form

OpenStudy (anonymous):

a function f(x,y) is said to be a homogeneouss de of degree n if f(ax,ay)=a^n f(x,y) for any non zero constant a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do you get f(ax,ay)=a^n f(x,y) here

OpenStudy (anonymous):

see i divided by x^2 here. everything's okay except for the 1/x^2.

OpenStudy (anonymous):

here its not a polynomial to have f(ax,ay)=a^n f(x,y) in differential eqautions its different

OpenStudy (anonymous):

Its not a homogeneous diff equation

OpenStudy (anonymous):

it came in todays foreign cbse board paper.